OpenStudy (tylerbanks2001):

Solve for w. 5w + 9z = 2z + 3w

OpenStudy (anonymous):

Move all terms containing w to the left, all other terms to the right. Add '-3w' to each side of the equation. 5w + -3w + 9z = 3w + -3w + 2z Combine like terms: 5w + -3w = 2w 2w + 9z = 3w + -3w + 2z Combine like terms: 3w + -3w = 0 2w + 9z = 0 + 2z 2w + 9z = 2z Add '-9z' to each side of the equation. 2w + 9z + -9z = 2z + -9z Combine like terms: 9z + -9z = 0 2w + 0 = 2z + -9z 2w = 2z + -9z Combine like terms: 2z + -9z = -7z 2w = -7z Divide each side by '2'. w = -3.5z Simplifying w = -3.5z

rebeccaxhawaii (rebeccaxhawaii):

add like terms